a square swimming pool with a side measure of 16 m is to be surrounded by a rubberized floor covering of uniform width. if the area of the floor covering equals the area of the pool, find the width of the rubberized covering.
Answer by texttutoring(324)

The area of the pool is A = 16*16 = 256
The area of the rubber floor is the same A=256

Let x = the width of the rubber covering.

Adding up the rubber area by breaking it up into 4 pieces (try drawing a picture to visualize it), we find:
16x + 16x + (16-2x) + (16-2x) = 256
32x + 32 - 4x = 256
28x = 224
x = 8

If you need a picture of it, let me know.